A(n) 11-inch sunflower is planted in a garden and the height of the sunflower increases exponentially. The height of the sunflower increases by 25% every 4 days.
What is the 4-day growth factor for the height of the sunflower?
What is the 1-day growth factor for the height of the sunflower?
What is the 1-day percent change for the height of the sunflower?
Write a function f that determines the height of the sunflower (in inches) in terms of the number of days since the sunflower was planted.

Found 2 solutions by Theo, josgarithmetic:
Answer by Theo(13342)

plant is 11 inches high to start with.
growth factor is 1.25 every 4 days.
answers to your questions are shown below:

(a) What is the 4-day growth factor for the height of the sunflower?

4 days growth factor is 1.25

(b) What is the 1-day growth factor for the height of the sunflower?

1 day growth factor is 1.25 ^ (1/4) = 1.057371263 rounded to the number of decimal digits that can be displayed.

(c) What is the 1-day percent change for the height of the sunflower?

1 day percent change for the height of the unflower is (1.057371263 - 1) * 100 = 5.7371263%.

((d) Write a function f that determines the height of the sunflower (in inches) in terms of the number of days since the sunflower was planted.

h = 11 * 1.25 ^ (x/4)

x is the number of days.

you can graph this function as shown below:



after 1 day, the plant has grown by a factor of 1.25 ^ (1/4).
11 * 1.25 ^ (1/4) is equal to 11.6310839.
the growth factor is 1.25 ^ (1/4) = 1.057371263.
the percent change is 5.7371263%.

by using the 4 day growth factor and dividing the number of days by 4, you automatically get the 1 day growth factor.



Answer by josgarithmetic(39623)

How is the "11 inch sunflower" meant? Say precisely!